In two years,Linda will be half as old as her mother. Her mother is 38 years old now. What is Linda's current age?

Answer by stanbon(75887)

In two years,Linda will be half as old as her mother. Her mother is 38 years old now. What is Linda's current age?
-----------
Equations:
L + 2 = (1/2)(M+2)
M = 38
-------------------
Substitute for "M" and solve for "L"::
L + 2 = (1/2)(38-2)
L + 2 = 18
L = 16 yrs (Linda's age now)
----------
Cheers,
Stan H.
--------